# LLMLingua-2: Data Distillation for Efficient and Faithful Task-Agnostic Prompt Compression

### v1

@@ 2 Related Works
Task-agnostic methods compress the prompt without considering the specific task, making it more adaptable to a range of applications and black-box LLMs. However, producing compressed text that can generalize well to different tasks is not trivial. Typical methods involve using information entropy-based metrics to remove redundant information in the prompt (Li et al., 2023; Jiang et al., 2023a). They employ a small language model to estimate token importance from the information metrics. Despite being training-free, these methods may not effectively capture the token importance distribution optimized for specific LLMs and often entail high computation overhead. Summarization-based methods are also leveraged for task-agnostic compression (Chen et al., 2023; Packer et al., 2023). However, they often omit crucial details and do not generalize well. An alternative approach is to compress or trim the context hidden or KV caches (Chevalier et al., 2023; Ge et al., 2023; Zhang et al., 2023; Liu et al., 2023b; Xiao et al., 2024). However, this is orthogonal to our work and cannot be easily applied to black-box LLMs.
@@ 3.1 Data Distillation
Our Instruction for Compression: Compress the given text to short expressions, and such that you (GPT-4) can reconstruct it as close as possible to the original. Unlike the usual text compression, I need you to comply with the 5 conditions below: 1. You can ONLY remove unimportant words. 2. Do not reorder the original words. 3. Do not change the original words. 4. Do not use abbreviations or emojis. 5. Do not add new words or symbols. Compress the origin aggressively by removing words only. Compress the origin as short as you can, while retaining as much information as possible. If you understand, please compress the following text: {text to compress} The compressed text is:
@@ 3.2 Data Annotation
Original Texts Item 15, report from City Manager Recommendation to adopt three resolutions. First, to join the Victory Pace program. Second, to join the California first program. And number three, consenting to to inclusion of certain properties within the jurisdiction in the California Hero program. Compressed Texts City Manager Recommendation adopt three resolutions. Join California first program. Consent properties inclusion jurisdiction California Hero program.
@@ 3.3 Quality Control
[…]
@@
[…]
@@
[…]
@@
F1 Score BELU Rouge1 Rouge2 RougeL BERTScore Tokens $1/\tau$
@@ 4.2 Compression Strategy
It’s worth noting that our approach can be readily integrated into the coarse-to-fine framework proposed in LLMLingua (Jiang et al., 2023a), allowing for a higher compression ratio of $\sim$15x for tasks involving multiple demonstrations or documents. Particularly, we can replace the perplexity-based iterative token compression module in LLMLingua with our token-classification-based compressor, while keeping the budget controller unchanged.
@@
${}^{{\dagger}}$ 33.8 35.9 25.9 23.5 18.0 17.8 25.8 1,947 5x 20.5 1,773 6x
${}^{{\dagger}}$ 34.3 36.3 24.7 32.4 26.3 24.8 29.8 1,991 5x 20.6 1,784 5x
${}^{{\dagger}}$ 39.0 42.2 27.4 69.3 53.8 56.6 48.0 1,809 6x 32.5 1,753 6x
@@
${}^{{\dagger}}$ 35.3 37.4 26.7 63.4 51.0 34.5 41.4 3,399 3x 24.0 3,340 3x
${}^{{\dagger}}$ 34.5 38.6 26.8 63.4 49.6 37.6 41.7 3,421 3x 22.4 3,362 3x
${}^{{\dagger}}$ 40.7 46.2 27.2 70.6 53.0 55.2 48.8 3,283 3x 32.8 3,412 3x
@@
Table 2: Out-of-domain evaluation on general long-context scenarios. ${}^{{\dagger}}$: numbers reported in Jiang et al. (2023b).
@@
Table 3: Out-of-domain evaluation on reasoning and in-context learning. ${}^{{\dagger}}$: numbers reported in Jiang et al. (2023b).
@@ 5 Experiment
(i) In-Domain: As we train our compressor using the dataset built with training examples from MeetingBank (Hu et al., 2023), we use the MeetingBank test examples for in-domain evaluation. In addition to the summarization task, we further introduce a QA task by prompting GPT-4 to generate 3 question-answer pairs for each example distributed across the whole context (see Appendix F for more details). For the summarization task, we use the same evaluation metric as in LLMLingua (Jiang et al., 2023a). For QA task, we use the metrics and scripts provided in LongBench (Bai et al., 2023) Single Document QA for evaluation.
@@
Table 6: Out-of-domain evaluation on general long-context benchmarks with the 2,000-token constraint. ${}^{{\ddagger}}$: We expand the constructed text compression dataset using 50k examples from TriviaQA-wiki. Then train an LLMLingua-2 compressor with the expanded dataset.
Instruction $1/\tau$ VR normal-$\downarrow$ QA F1 normal-$\uparrow$
@@ 6 Conclusion
Secondly, we expand the constructed text compression dataset using 50k examples from TriviaQA-wiki. Then train an LLMLingua-2 compressor with the expanded dataset to see whether there would be further performance gain. Table 6 shows the results under the 2,000-token constraint. We can see that training the compressor with more data does bring further performance gain (${}^{{\ddagger}}$). However, the improvement seems not that significant. We conjecture that this is because although the semantics of texts from different domains may vary a lot, their redundancy pattern might be similar. Such pattern or knowledge may be learned during in-domain training, and then act as an anchor that can transfer across different domains. We leave this for future work.
@@ Appendix B Details of Data Annotation
Another challenge arises from the fact that GPT-4 may alter the original words in tense, voice, and singular/plural forms during compression, even when we request GPT-4 to compress by discarding words only. To address this issue, we first apply lemmatization to reduce words to their base form using Spacy [^5] https://spacy.io/api/lemmatizer, and then perform word matching using the sliding window approach.
@@ Appendix E Prompt Reconstruction
Our GPT-4 Instruction for Compression: System Prompt: You are an excellent linguist and very good at compressing passages into short expressions by removing unimportant words, while retaining as much information as possible. User Prompt: Compress the given text to short expressions, and such that you (GPT-4) can reconstruct it as close as possible to the original. Unlike the usual text compression, I need you to comply with the 5 conditions below: 1. You can ONLY remove unimportant words. 2. Do not reorder the original words. 3. Do not change the original words. 4. Do not use abbreviations or emojis. 5. Do not add new words or symbols. Compress the origin aggressively by removing words only. Compress the origin as short as you can, while retaining as much information as possible. If you understand, please compress the following text: {text to compress} The compressed text is:
@@
Instruction1: Could you please rephrase the paragraph to make it short, and keep 5% tokens? Instruction2: Summarize the provided examples in a few sentences, maintaining all essential reasoning aspects. Instruction3: Remove redundancy and express the text concisely in English, ensuring that all key information and reasoning processes are preserved. Instruction4: Follow these steps to shorten the given text content: 1. First, calculate the amount of information contained in each sentence, and remove sentences with less information. 2. Next, further condense the text by removing stop words, unnecessary punctuation, and redundant expressions. Refine the content while ensuring that all key information is retained. Let’s do it step by step.
@@ Appendix I GPU Memory Usage
As illustrated in Table 11, LLMLingua-2 with LongLLMLingua coarse-grained compression achieves an average performance gain of 25.3% on NaturalQuestions (Liu et al., 2023a) compared to LLMLingua-2.
@@
LLMLingua 35.2 20.4 11.8 24.3 51.4 28.6 3060 5x
LLMLingua-2 46.7 23.0 15.3 32.8 72.6 38.1 3023 5x
Original Prompt 61.2 28.7 16.0 29.2 77.5 42.5 14940 -
@@
Data Type QA Summary Length
F1 Score BELU Rouge1 Rouge2 RougeL BERTScore # Tokens $1/\tau$
Annotated Filtered 58.71 17.74 48.42 23.71 34.36 88.99 1629 3.3x
Kept 92.82 19.53 50.24 25.16 36.38 89.05 855 2.9x
All 86.30 19.17 49.89 24.90 35.97 89.04 1003 3.0x
Original Filtered 59.65 20.53 46.39 25.31 34.17 88.91 5298 -
Kept 94.41 23.05 47.73 27.20 35.74 88.99 2461 -
All 87.75 22.34 47.28 26.66 35.15 88.96 3,003 -
Table 10: Ablation study of the filtering process in dataset construction. Annotated gathers all words which are assigned a True label by our annotation algorithm as the input prompt. Filtered denotes the discard samples of the filtering process in sec 3.3, while Kept represents the retained samples.
@@
${}^{{\dagger}}$ 40.6 38.6 38.2 37.4 36.6 36.3 798 3.7x
${}^{{\dagger}}$ 63.1 61.0 59.8 61.1 60.1 62.3 824 3.6x
${}^{{\dagger}}$ 66.9 61.1 59.0 61.2 60.3 64.4 808 3.6x
${}^{{\dagger}}$ 63.8 64.6 65.4 64.1 63.7 63.7 804 3.7x
${}^{+}$ 74.0 70.4 67.0 66.9 65.3 71.9 739 3.9x
${}^{{\dagger}}$ 75.0 71.8 71.2 71.2 74.7 75.5 748 3.9x
@@
${}^{{\dagger}}$ 31.4 19.5 24.7 24.1 43.8 - 791 3.7x
${}^{{\dagger}}$ 25.5 27.5 23.5 26.5 30.0 27.0 775 3.8x
LLMLingua2 48.6 44.5 43.6 40.9 39.9 46.2 748 3.9x
@@
Table 11: Performance comparison on NaturalQuestions (20 documents) (Liu et al., 2023a). ${}^{+}$ denotes LLMLingua-2 with LongLLMLingua Jiang et al. (2023b) coarse level compression. ${}^{{\dagger}}$: numbers reported in Jiang et al. (2023b).
@@
Target Token Constraint 2000 Tokens 3000 Tokens
LLMLingua2 29.8 1954 7.4x 35.5 3392 4.3x
@@
LLMLingua2 ${}^{{\dagger}}$ 25.1 2131 6.8x 27.4 3185 4.5x
LLMLingua2 ${}^{{\ddagger}}$ 29.5 2125 6.8x 32.2 3164 4.5x
@@
Table 12: Evaluation of LLMLingua-2 sample wise dynamic compression on LongBench single doc QA task. ${}^{{\dagger}}$ assigns each example with the same fixed compression rate. ${}^{{\ddagger}}$ assigns dynamic compression rate to different examples within the corpus level constraint.

### v2

@@ 2 Related Works
Task-agnostic methods compress the prompt without considering the specific task, making it more adaptable to a range of applications and black-box LLMs. However, producing compressed text that can generalize well to different tasks is not trivial. Typical methods involve using information entropy-based metrics to remove redundant information in the prompt (Li et al., 2023; Jiang et al., 2023a). They employ a small language model to estimate token importance from the information metrics. Despite being training-free, these methods may not effectively capture the token importance distribution optimized for specific LLMs and often entail high computation overhead. Summarization-based methods are also leveraged for task-agnostic compression (Chen et al., 2023; Packer et al., 2023). However, they often omit crucial details and do not generalize well. An alternative approach is to compress or trim the context hidden or KV caches (Chevalier et al., 2023; Ge et al., 2024; Zhang et al., 2023; Liu et al., 2023; Xiao et al., 2024). However, this is orthogonal to our work and cannot be easily applied to black-box LLMs.
@@ 3.1 Data Distillation
Our Instruction for Compression:
Compress the given text to short expressions, and such that you (GPT-4) can reconstruct it as close as possible to the original. Unlike the usual text compression, I need you to comply with the 5 conditions below:
1. You can ONLY remove unimportant words. 2. Do not reorder the original words. 3. Do not change the original words. 4. Do not use abbreviations or emojis. 5. Do not add new words or symbols. Compress the origin aggressively by removing words only. Compress the origin as short as you can, while retaining as much information as possible. If you understand, please compress the following text: {text to compress} The compressed text is:
@@ 3.2 Data Annotation
Original Texts
Item 15, report from City Manager Recommendation to adopt three resolutions. First, to join the Victory Pace program. Second, to join the California first program. And number three, consenting to to inclusion of certain properties within the jurisdiction in the California Hero program.
Compressed Texts
City Manager Recommendation adopt three resolutions. Join California first program. Consent properties inclusion jurisdiction California Hero program.
@@ 3.3 Quality Control
$\textit{VR}=\frac{1}{|\mathbb{S}_{comp}|}\sum_{w\in\mathbb{S}_{comp}}\mathbb{I}(w\notin\mathbb{S}_{ori}),$ (1)
@@
We propose Alignment Gap (AG) to evaluate the quality of the automatically annotated labels. Let $l(\cdot)$ represent the annotation function, where $l(w)=\textit{True}$ signifies that word $w\in\mathbb{S}_{ori}$ corresponds to a word in $\mathbb{S}_{comp}$. We firstly define the matching rate (MR) as:
$\textit{MR}=\frac{1}{|\mathbb{S}_{ori}|}\sum_{w\in\mathbb{S}_{ori}}\mathbb{I}(l(w)=\textit{True}).$ (2)
@@
$\textit{HR}=\frac{1}{|\mathbb{S}_{ori}|}\sum_{w\in\mathbb{S}_{comp}}\mathbb{I}(w\in\mathbb{S}_{ori}).$ (3)
@@
EM BLEU Rouge1 Rouge2 RougeL BERTScore Tokens $1/\tau$
@@ 4.2 Compression Strategy
It’s worth noting that our approach can be readily integrated into the coarse-to-fine framework proposed in LLMLingua (Jiang et al., 2023a), allowing for a higher compression ratio of $\sim$15x for tasks involving multiple demonstrations or documents. Particularly, we can replace the perplexity-based iterative token compression module in LLMLingua with our token-classification-based compressor, while keeping the budget controller unchanged. Detailed information can be found in Appendix K.
@@
SBERT 33.8 35.9 25.9 23.5 18.0 17.8 25.8 1,947 5x 20.5 1,773 6x
OpenAI 34.3 36.3 24.7 32.4 26.3 24.8 29.8 1,991 5x 20.6 1,784 5x
LongLLMLingua 39.0 42.2 27.4 69.3 53.8 56.6 48.0 1,809 6x 32.5 1,753 6x
@@
SBERT 35.3 37.4 26.7 63.4 51.0 34.5 41.4 3,399 3x 24.0 3,340 3x
OpenAI 34.5 38.6 26.8 63.4 49.6 37.6 41.7 3,421 3x 22.4 3,362 3x
LongLLMLingua 40.7 46.2 27.2 70.6 53.0 55.2 48.8 3,283 3x 32.8 3,412 3x
@@
Table 2: Out-of-domain evaluation on general long-context scenarios. : numbers reported in Jiang et al. (2023b).
@@
Table 3: Out-of-domain evaluation on reasoning and in-context learning. : numbers reported in Jiang et al. (2023b).
@@ 5 Experiment
(i) In-Domain: As we train our compressor using the dataset built with training examples from MeetingBank (Hu et al., 2023), we use the MeetingBank test examples for in-domain evaluation. In addition to the summarization task, we further introduce a QA task by prompting GPT-4 to generate 3 question-answer pairs for each example distributed across the whole context (see Appendix F for more details). For the summarization task, we use the same evaluation metric as in LLMLingua (Jiang et al., 2023a). For QA task, we utilize the Exact Match as the evaluation metric.
@@
Table 6: Out-of-domain evaluation on general long-context benchmarks with the 2,000-token constraint. LLMLingua-2: We expand the constructed text compression dataset using 50k examples from TriviaQA-wiki. Then train an LLMLingua-2 compressor with the expanded dataset.
Instruction $1/\tau$ VR $\downarrow$ QA F1 $\uparrow$
@@ 6 Conclusion
Secondly, we expand the constructed text compression dataset using 50k examples from TriviaQA-wiki. Then train an LLMLingua-2 compressor with the expanded dataset to see whether there would be further performance gain. Table 6 shows the results under the 2,000-token constraint. We can see that training the compressor with more data does bring further performance gain (LLMLingua-2). However, the improvement seems not that significant. We conjecture that this is because although the semantics of texts from different domains may vary a lot, their redundancy pattern might be similar. Such pattern or knowledge may be learned during in-domain training, and then act as an anchor that can transfer across different domains. We leave this for future work.
@@ Appendix B Details of Data Annotation
Another challenge arises from the “variation" misbehavior of GPT-4, as illustrated in Fig 5. GPT-4 may alter the original words in tense, voice, and singular/plural forms during compression, even when we request GPT-4 to compress by discarding words only. To address this issue, we first apply lemmatization to reduce words to their base form using Spacy [^5] https://spacy.io/api/lemmatizer, and then perform word matching using the sliding window approach.
@@ Appendix E Prompt Reconstruction
Our GPT-4 Instruction for Compression: System Prompt:
You are an excellent linguist and very good at compressing passages into short expressions by removing unimportant words, while retaining as much information as possible.
User Prompt:
Compress the given text to short expressions, and such that you (GPT-4) can reconstruct it as close as possible to the original. Unlike the usual text compression, I need you to comply with the 5 conditions below:
1. You can ONLY remove unimportant words. 2. Do not reorder the original words. 3. Do not change the original words. 4. Do not use abbreviations or emojis. 5. Do not add new words or symbols. Compress the origin aggressively by removing words only. Compress the origin as short as you can, while retaining as much information as possible. If you understand, please compress the following text: {text to compress} The compressed text is:
@@
Instruction1:
Could you please rephrase the paragraph to make it short, and keep 5% tokens?
Instruction2: Summarize the provided examples in a few sentences, maintaining all essential reasoning aspects. Instruction3: Remove redundancy and express the text concisely in English, ensuring that all key information and reasoning processes are preserved. Instruction4: Follow these steps to shorten the given text content: 1. First, calculate the amount of information contained in each sentence, and remove sentences with less information. 2. Next, further condense the text by removing stop words, unnecessary punctuation, and redundant expressions. Refine the content while ensuring that all key information is retained. Let’s do it step by step.
@@ Appendix I GPU Memory Usage
As illustrated in Table 11, LLMLingua-2 with LongLLMLingua coarse-grained compression achieves an average performance gain of 25.3% on NaturalQuestions (Liu et al., 2024) compared to LLMLingua-2.
@@
Appendix M Performance w.r.t Compression Ratio
Fig 15 presents the performance w.r.t compression ratio on a subset of 100 samples from Meetingbank. As depicted, LLMLingua-2 exhibits superior robustness compared to other baselines as the compression ratio increases.
Appendix N Preservation Priority in GPT-4 Compression
To gain insight into GPT-4’s compression patterns, we analyze the distribution of different POS categories. Fig 16 suggests that GPT-4 prioritizes the preservation of nouns, adjectives, and numerals, which typically play a more important role in the comprehension of the overall context.
Appendix O Comparison With GPT-4 Compression
Table 9 shows the comparison between LLMLingua-2 compressed prompts and GPT-4 compressed prompts. For GPT-4 compression, We use the same compression instruction as the one used in training data collection. The same chunking technique is also adopted with the chunk size setting to 512. It is shown that LLMLingua-2 achieves higher performance than GPT-4 compression on MeetingBank QA. We conjecture that LLMLingua-2’s ability to learn compression knowledge from the entire dataset helps mitigate the influence of noise and information loss present in each GPT-4 compressed example, leading to superior performance.
Methods QA Length
EM Tokens $1/\tau$
GPT-4 Compression 84.86 1,221 2.5x
LLMLingua-2-small 85.82 984 3.0x
LLMLingua-2 86.92 970 3.1x
Original 87.75 3,003 1.0x
Table 9: Comparison with GPT-4 compressed prompt on MeetingBank.
Appendix P Performance of Mistral-7B on 8K Token Subset
As the Mistral 7B model is trained with an 8k context length  [^6] https://huggingface.co/docs/transformers/main/en/model_ doc/mistral , its performance may drop if the input context is too long. Therefore, we conduct additional experiments on subsets containing only examples with original prompts shorter than 8k tokens. The results, shown in Table 13, demonstrate that LLMLingua-2 continues to outperform strong baselines and even the original prompts in this subset.
@@
LLMLingua 35.2 20.4 11.8 24.3 51.4 28.6 3,060 5x
LLMLingua-2 46.7 23.0 15.3 32.8 72.6 38.1 3,023 5x
Original Prompt 61.2 28.7 16.0 29.2 77.5 42.5 14,940 -
@@
BM25 40.6 38.6 38.2 37.4 36.6 36.3 798 3.7x
Gzip 63.1 61.0 59.8 61.1 60.1 62.3 824 3.6x
SBERT 66.9 61.1 59.0 61.2 60.3 64.4 808 3.6x
OpenAI 63.8 64.6 65.4 64.1 63.7 63.7 804 3.7x
LLMLingua-2+ 74.0 70.4 67.0 66.9 65.3 71.9 739 3.9x
LongLLMLingua 75.0 71.8 71.2 71.2 74.7 75.5 748 3.9x
@@
Selective-Context 31.4 19.5 24.7 24.1 43.8 - 791 3.7x
LLMLingua 25.5 27.5 23.5 26.5 30.0 27.0 775 3.8x
LLMLingua-2 48.6 44.5 43.6 40.9 39.9 46.2 748 3.9x
@@
Table 11: Performance comparison on NaturalQuestions (20 documents) (Liu et al., 2024). LLMLingua-2+ denotes LLMLingua-2 with LongLLMLingua Jiang et al. (2023b) coarse level compression. : numbers reported in Jiang et al. (2023b).
@@
Target Token Constraint 2,000 Tokens 3,000 Tokens
LLMLingua-2 29.8 1,954 7.4x 35.5 3,392 4.3x
@@
LLMLingua-2 FR 25.1 2,131 6.8x 27.4 3,185 4.5x
LLMLingua-2 DCR 29.5 2,125 6.8x 32.2 3,164 4.5x
@@
Table 12: Evaluation of LLMLingua-2 sample wise dynamic compression on LongBench single doc QA task. FR assigns each example with the same fixed compression rate. DCR assigns dynamic compression rate to different examples within the corpus level constraint.
(a) QA performance w.r.t compression ratio on a 100 samples subset of MeetingBank.
(b) Summary performance w.r.t compression ratio on a 100 samples subset of MeetingBank.
Figure 15: A plot of performance w.r.t compression ratio on a 100 samples subset of MeetingBank.
Figure 16: Part of speech distribution of the original prompts and GPT-4 compressed prompts.
Methods MeetingBank LongBench-SingleDoc
Selective-Context 62.43 19.25 703 2.4x 29.3 1,829 2.5x 34.6 2,855 1.6x
LLMLingua 51.78 24.57 714 2.4x 29.9 1,862 2.5x 30.7 3,016 1.5x
LLMLingua-2 81.75 30.83 651 2.6x 35.0 1,889 2.4x 36.3 2,841 1.6x
Original Prompt 71.27 27.56 1,700 - 31.4 4,595 - 31.4 4,595 -
Table 13: Evaluation with Mistral-7B as the Target LLM on MeetingBank and LongBench single doc QA task. We discarded samples where the input text has more than 8K tokens. We report Rouge1Lin (2004) for summary.